OpenStudy (anonymous):

Solve by elimination: 5x+3y=-4 x+y=-4

OpenStudy (ksaimouli):

add them

OpenStudy (anonymous):

ummmm that would be 8xy but I don't think that's right... because... I don't remember learning that...

OpenStudy (anonymous):

then you I don't know what to do with the -4

OpenStudy (anonymous):

\[5x+3y=-4; x+y=-4 ;\] multiply eq 2 with 5\[5x+5y=-20.\] substract from eq 1.\[5x+3y-5x-5y=-4+20; -2y=16 ; y= -8 \] put y=-8 in either one of the eq. x=4